OpenStudy (anonymous):

The sum of two numbers is 46. The sum of the smaller and 3 times the larger is 104. Find the numbers.

OpenStudy (callisto):

let x and y be the smaller and larger no respectively x+y = 46 => x=46-y (1) x+3y = 104 (2) Put (1) into (2) y = 29, x=46-29=17

OpenStudy (anonymous):

let smaller no = x larger no = y x + y = 46 -------- i x + 3y = 104 ------- ii subtract ii from i -2y = - 58 y = -58/-2 y = 29 so larger no is 29 smaller no is 46 - 29 = 17

OpenStudy (anonymous):

Thank you!